### v1

# Crossing the Human-Robot Embodiment Gap with Sim-to-Real RL using One Human Demonstration

Abstract
Teaching robots dexterous manipulation skills often requires collecting hundreds of demonstrations using wearables or teleoperation, a process that is challenging to scale. Videos of human-object interactions are easier to collect and scale, but leveraging them directly for robot learning is difficult due to the lack of explicit action labels from videos and morphological differences between robot and human hands. We propose Human2Sim2Robot, a novel real-to-sim-to-real framework for training dexterous manipulation policies using only one RGB-D video of a human demonstrating a task. Our method utilizes reinforcement learning (RL) in simulation to cross the human-robot embodiment gap without relying on wearables, teleoperation, or large-scale data collection typically necessary for imitation learning methods. From the demonstration, we extract two task-specific components: (1) the object pose trajectory to define an object-centric, embodiment-agnostic reward function, and (2) the pre-manipulation hand pose to initialize and guide exploration during RL training. We found that these two components are highly effective for learning the desired task, eliminating the need for task-specific reward shaping and tuning. We demonstrate that Human2Sim2Robot outperforms object-aware open-loop trajectory replay by 55% and imitation learning with data augmentation by 68% across grasping, non-prehensile manipulation, and multi-step tasks. Project Site: human2sim2robot.github.io
I Introduction
Training robotic manipulation policies from demonstrations via imitation learning (IL) has gathered considerable interest [1, 2, 12, 19, 23]. A key requirement for effective IL is the availability of high-quality demonstrations composed of sequences of states with corresponding robot action labels. This is typically acquired through labor-intensive data collection requiring hundreds or thousands of human demonstrations per task via teleoperation or specialized wearable equipment [10, 28, 53]. In contrast, videos of humans interacting with objects with their own hands are inexpensive to collect, a scalable alternative to traditional demonstration collection methods. However, leveraging human videos directly for IL is challenging as they lack explicit robot action labels [57].
Figure 1: Human2Sim2Robot is a real-to-sim-to-real RL framework for learning dexterous manipulation policies from one human RGB-D video demonstration per task. From the demonstration, we leverage object pose trajectories and pre-manipulation hand poses to provide guidance and advantageous initialization for RL policy training in simulation. These policies demonstrate zero-shot sim-to-real transfer to a Kuka arm and Allegro hand.
One popular approach to address this is by obtaining per-timestep human hand pose estimates and converting them to robot action labels via fingertip retargeting and inverse kinematics (IK) [4, 53]. However, this approach faces two challenges. First, while recent hand pose reconstruction methods have shown impressive performance [38], their predictions are still not always accurate due to occlusions and sensor noise, resulting in mismatches between the robot hand IK targets and the true human hand fingertips. Second, even with perfect human hand pose estimates, the considerable embodiment gap between human and robot hands may make solving IK infeasible or result in suboptimal robot joint configurations for the demonstrated task. Both challenges are particularly punishing for contact-rich, dexterous manipulation tasks using multi-fingered robotic hands as they require precise coordination of forces and motion at multiple contact points, which is difficult to achieve when retargeted actions are noisy or suboptimal for the target embodiment [3, 32, 37]. Existing IL methods are ill-equipped to address this issue as they rely on accurate correspondences between the demonstrated and learned behaviors. Reinforcement learning (RL) offers a promising alternative to overcome these limitations by enabling the robot to directly learn manipulation tasks using its own embodiment. However, RL has several limitations such as tedious, task-specific reward engineering and unfavorable sample complexity, which make training real-world policies infeasible [49, 60].
We propose Human2Sim2Robot, a real-to-sim-to-real RL framework that addresses these limitations by learning robust dexterous manipulation policies from a single human hand RGB-D video demonstration. We extract two task-specific components from the video: (1) the object 6D pose trajectory, which defines an object-centric reward function that specifies the desired task while accommodating embodiment differences, and (2) the pre-manipulation hand pose, which provides advantageous initialization of RL training such that the resulting policy converges on a similar manipulation strategy shown in the demonstration. Formalizing an expert task demonstration with these two components facilitates RL policy training with no task-specific reward tuning. Instead of directly learning state-action mappings, we use the demonstration for task specification and guidance, encouraging human-like behavior when beneficial while allowing deviations when the human strategy is unsuitable for the robot’s embodiment. Human2Sim2Robot policies achieve zero-shot sim-to-real transfer on a real dexterous robot, without requiring wearable technologies, teleoperation, or large-scale data collection.
To the best of our knowledge, Human2Sim2Robot is the first system that learns a robust real-world dexterous manipulation policy from only one human RGB-D video demonstration, bridging the human-robot embodiment gap for a wide range of grasping, non-prehensile manipulation, and multi-step tasks. Our extensive ablation studies demonstrate the importance of our system’s design decisions. While individual components have precedents, these works are often limited to simulation [42, 47], are not reactive closed-loop policies [6, 22, 56], require significantly more demonstrations [25, 42, 47, 58], or only perform prehensile manipulation [42, 47, 56, 22]. We demonstrate that Human2Sim2Robot is highly effective for training task-specific RL policies in simulation. The resulting policies can execute diverse real-world dexterous manipulation tasks, such as pouring from a pitcher, pivoting a box against a wall, inserting a plate into a dishrack, or even composed sequences of tasks, without any task-specific reward tuning. In the challenging single demonstration regime, Human2Sim2Robot outperforms object-aware open-loop trajectory replay by 55% and imitation learning with data augmentation by 68% across grasping, non-prehensile and extrinsic manipulation, and complex multi-step tasks.
II Related Work
II-A Imitation Learning for Robotics
II-A1 Visuomotor Imitation Learning
Recent work in visuomotor IL for robotic manipulation has shown success learning from a large number of expert demonstrations [1, 2, 9, 19, 23, 28, 53, 59]. Demonstrations are typically collected through teleoperation or specialized wearable equipment [10, 28, 53] like motion capture gloves, AR/VR equipment, or portable robot hands, which makes scaling of data collection efforts expensive. In contrast, videos of human demonstrations are inexpensive to collect and more intuitive to demonstrators. Since videos lack explicit action labels, a popular approach is to obtain per-timestep human hand pose estimates and convert them into robot action labels through IK-based retargeting [4, 53]. However, the human-robot embodiment gap often makes finding an IK solution infeasible or result in a robot joint configuration that is suboptimal for reproducing the demonstrated task. This poses significant challenges for visuomotor IL methods that directly rely on accurate correspondences between the demonstrated and learned behaviors.
Therefore, instead of directly learning a mapping from observations to actions, Human2Sim2Robot performs RL in simulation guided by a single human video demonstration. This approach acknowledges that while human demonstrations provide useful guiding strategies for completing the task, certain actions may not be suitable for robots given substantial embodiment differences. Learning dexterous manipulation policies with RL hence encourages human-like behavior when beneficial while allowing deviations when the human strategy is unsuitable for the robot’s embodiment. Furthermore, robust recovery and retry behavior emerges as a result of training and does not need to be explicitly demonstrated as in IL.
II-A2 One-Shot & Few-Shot Imitation Learning
Another category of IL methods that parallels our approach follows the one-shot IL (OSIL) paradigm, where a single demonstration is provided. Past work has performed object pose estimation [18, 51] or object-aware retargeting using open-world vision models [27] to transfer the demonstrated trajectory to novel scenes, or adapted the single demonstration to a new scene by leveraging object segmentation and visual servoing to move the robot’s end effector to the same position relative to the object before replaying the demonstration [50]. Beyond strictly using one demonstration, a related approach collects a small number ($\sim$5) of teleoperated demonstrations per task, then generates more data by retargeting demonstrations to new initial conditions and rolling out these trajectories in a simulated digital twin, executing these actions in the real world if they successfully complete the task in simulation [20].
While these approaches are more data-efficient than visuomotor IL policies, they suffer from limited generalization beyond the demonstrated actions. Furthermore, when learning from human video demonstrations, object and hand pose estimates are often inaccurate due to occlusions and noise. Even if pose estimates are accurate, the human-robot embodiment gap causes IK-based retargeting to produce infeasible robot hand trajectories. Therefore, simply replaying modified versions of the single demonstration or directly learning observation-to-action mappings from retargeted data is unlikely to succeed. Our insight is that using the single human video demonstration to provide task specification and guidance for RL more effectively leverages this data source for robot learning. This allows robots to develop effective strategies with their own embodiment, rather than rigidly imitating human behaviors.
II-B Reinforcement Learning for Robotics
Reinforcement learning (RL) is a method for training autonomous agents to perform complex tasks by interacting with an environment through trial and error. In this work, we train manipulation policies in simulation to avoid the pitfalls of RL in the real world such as slow training, unsafe behavior, frequent environment resets and difficult to tune reward functions [49, 60]. Sim-to-real RL has shown significant promise across a wide range of robotic tasks, achieving state-of-the-art performance in domains such as legged locomotion [7, 35, 36], drone racing [21], bipedal soccer [17], and in-hand manipulation [41]. However, this potential has yet to be fully realized for dexterous manipulation over the full robot arm-and-hand kinematics: many prior works on RL for dexterous manipulation are confined to simulation with non-physical, floating-hand robots [5, 43, 52, 55].
Recent works leveraging sim-to-real RL for real-world manipulation tasks have used RL to fine-tune a BC policy [49] or learn a residual policy that outputs delta actions relative to an open-loop base motion [6]. Fine-tuning a BC policy may learn more stably, but requires a demonstration dataset with the same robot embodiment. Residual policy learning can be effective, but only allows small adjustments to the base motion, limiting its ability to overcome large embodiment gaps or demonstrate retry behavior. Additionally, both methods require accurate human action sequences from high-quality motion capture or teleoperation. Human2Sim2Robot is a real-to-sim-to-real framework that trains an RL policy over a full arm-and-hand action space, guided by object-centric rewards and a pre-manipulation hand pose from just one human demonstration. This approach can thus accommodate lower-quality human video demonstration data, facilitate learning of retry behavior, and overcome the human-robot embodiment gap. Furthermore, our approach enables policies to learn contact-rich dexterous manipulation tasks that are more complex than parallel-jaw pick-and-place or in-hand manipulation with a static arm.
Lum et al. [30] trains dexterous grasping policies using sim-to-real RL with a geometric fabric controller, enabling smooth, coordinated motion while effectively avoiding undesired collisions with the environment. However, the policy is limited to a simple grasping task, typically converging on a simple top-down grasp as the hand is initialized above the object and the policy is trained from scratch without human guidance. Human2Sim2Robot leverages a similar geometric fabric controller, but enables policies to perform more general prehensile and non-prehensile manipulation tasks by incorporating human guidance on how to perform the task.
Recent studies corroborate the observation that pre-grasp hand configurations can accelerate policy learning and result in human-like grasps [11, 14, 31]. These methods only focus on grasping tasks, use pre-grasps from very similar embodiments, and have only been tested in simulation. In this work, we focus on training RL policies that transfer to the real-world, overcome the human-robot embodiment gap, and perform both prehensile and non-prehensile manipulation. Furthermore, we incorporate scalable methods for pre-manipulation hand pose acquisition from one human video demonstration.
Another line of work uses inverse RL on human videos [25, 58], but inferring a reward function typically requires $\sim$100 demos. In contrast, our explicit object-centric reward is designed for sample efficiency in the single-demo regime.
III Method
We present Human2Sim2Robot, a real-to-sim-to-real RL framework for learning robust, dexterous manipulation policies from a single human-hand RGB-D video demonstration. Figure 1 shows an overview of our framework. The following sections detail key design decisions of our framework.
III-A Real-to-Sim & Human Demonstration
We first create a digital twin of the robot’s real-world environment and the target object to act as a policy training ground in simulation. Concretely, we use off-the-shelf 3D LiDAR scanning apps, Kiri Engine [24] and 3D Scanner App [26], to capture a high-fidelity object mesh $\mathcal{O}$ and scene mesh $\mathcal{S}$, respectively, with a few minutes of human effort.
Next, we use a camera with known intrinsic and extrinsic parameters to record a single monocular RGB-D video demonstration $\{\bm{I}_{t}\}_{t=1}^{T}$, where $\bm{I}_{t}\in\mathbb{R}^{H\times W\times 4}$ is the RGB-D image at timestep $t\in\mathbb{N}$, $T\in\mathbb{N}$ is the total number of timesteps, and $H$ and $W$ are the image height and width, respectively.
Figure 2: Human Demonstration Processing. We extract two task-specific components from the human hand RGB-D video demonstration: (1) the object pose trajectory, which defines an object-centric, embodiment-agnostic reward function, and (2) the pre-manipulation hand pose, which provides advantageous initialization for RL policy training. We use FoundationPose [54] for object pose estimation (requires per-frame object masks obtained via SAM 2 [44] and object mesh [24]) and HaMeR [38] for hand pose estimation.
Figure 2 visualizes how we process the human demonstration to obtain (1) an object pose goal trajectory $\{\bm{T}^{\text{goal}}_{t}\}_{t=1}^{T}$, and (2) a human hand pose trajectory represented as MANO [45] parameters $\left\{(\bm{\theta}_{t},\bm{\beta}_{t})\right\}_{t=1}^{T}$. At each timestep $t$, $\bm{T}^{\text{goal}}_{t}\in\text{SE}(3)$ is the object pose from the demonstration’s object trajectory, $\bm{\theta_{t}}\in\mathbb{R}^{48}$ is the MANO hand pose parameter, and $\bm{\beta_{t}}\in\mathbb{R}^{10}$ is the MANO hand shape parameter. In our system, we extract the object pose trajectory using FoundationPose [54], an open-source object pose detection model, which requires the scanned object mesh $\mathcal{O}$ and per-timestep object masks generated using Segment Anything Model 2 (SAM 2) [44], an open-source segmentation model. We extract per-timestep human hand poses using HaMeR [38], an open-source hand pose detection model. Since HaMeR takes RGB images as input, we use accurate depth values from depth images and perform ICP registration to align the hand point clouds for obtaining accurate hand poses (see Appendix -A for details).
Next, we determine the pre-manipulation hand pose at timestep $\tau=t_{0}-t_{\text{offset}}$, where we define $t_{0}$ as the first timestep in which the object’s velocity exceeds a threshold $v_{\text{min}}=5cm/s$, and $t_{\text{offset}}=10$ represents a fixed number of timesteps prior to the object’s motion. We then use $\bm{\theta}_{\tau}$ and $\bm{\beta}_{\tau}$ to compute the fingertip positions and the middle finger base knuckle pose as the human pre-manipulation hand pose.
Lastly, we retarget the human pre-manipulation hand pose to a robot hand pose through a two-step IK procedure using cuRobo [48]. Figure 3 visualizes how we perform human-to-robot hand retargeting. In the first step, the robot arm’s joint angles are adjusted to align the base position and orientation of the robot hand’s middle knuckle with that of the human hand (with a small offset, see Appendix -B for details). In the second step, the robot hand’s joint angles are adjusted to align the robot’s fingertip positions with the corresponding human fingertip positions. This generates a pre-manipulation robot hand pose $(\bm{T}^{\text{wrist}},\bm{q}^{\text{hand}})$ for the object pose represented as $\bm{T}^{\text{goal}}_{\tau}$, where $\bm{T}^{\text{wrist}}\in\text{SE}(3)$ is the robot wrist pose, and $\bm{q}^{\text{hand}}\in\mathbb{R}^{N^{\text{hand-joints}}}$ is robot hand joint configuration. $N^{\text{hand-joints}}=23$ for the Allegro hand. This simple IK process faithfully retargets the human pre-manipulation hand pose, while maintaining kinematic feasibility and alignment between human and robot hand configurations.
The object pose trajectory is used for task specification by defining an object-centric trajectory-tracking reward function used for policy training. The human pre-manipulation hand pose is used for task guidance by defining a good reset state initialization for exploring how to perform the task [14]. We note that the pre-manipulation hand pose retargeting process does not need to be extremely precise, as it is only used as a prior from which to explore and improve from. Together, these abstractions jointly guide RL policy training in simulation via reward guidance and advantageous state initializations.
Figure 3: Human to Robot Hand Retargeting. (a) MANO hand pose estimated by HaMeR [38] and aligned using ICP registration. Colored spheres denote IK retargeting points on the human hand: middle knuckle (red) and fingertips (pink, green, blue, yellow). (b) IK Step 1 (Arm): Robot hand’s middle finger base is aligned with the human hand’s middle finger knuckle. (c) IK Step 2 (Hand): Fingertips of human and robot hands are aligned.
III-B Simulation-based Policy Learning
We create a training environment in the IsaacGym simulator [34] that matches the real-world environment, consisting of the dexterous robot, the scene mesh $\mathcal{S}$, and the object mesh $\mathcal{O}$. We then train a policy using Proximal Policy Optimization [46] which outputs robot actions that move the object along the desired goal trajectory, guided by the provided pre-manipulation hand pose. We emphasize that we primarily care about how the object moves, rather than imitating the actions of the human demonstration; the pre-manipulation hand pose serves as a rough prior, but the policy will learn to use the robot embodiment to achieve the desired object motion.
The reward given at timestep $t$ is a simple object-tracking reward, $r_{t}=r^{\text{obj}}_{t}$, defined as
$r^{\text{obj}}_{t}=\exp\bigl{(}-\alpha\,d(\bm{T}^{\text{goal}}_{\tau+t},\bm{T}^{\text{obj}}_{t})\bigr{)}$ (1)
We define the relative pose distance function, $d$, as
$\centering d(\bm{T}_{1},\bm{T}_{2})=\sum_{i=1}^{N^{\text{anchor}}}\bigl{\lVert}\,\bm{T}_{1}\bm{k}_{i}-\bm{T}_{2}\bm{k}_{i}\bigr{\rVert}\@add@centering$ (2)
For most objects, we select $N^{\text{anchor}}=3$, with $\bm{k}_{1}=[L,0,0]$, $\bm{k}_{2}=[0,L,0]$, and $\bm{k}_{3}=[0,0,L]$ in the local object frame, where $L$ is a distance parameter for orientation. Figure 4 visualizes the object pose trajectory tracking reward. This formulation integrates both position and orientation in a physically interpretable manner: larger values of $L$ increase the influence of orientation, since the anchor points lie farther from the object’s origin. Moreover, it is flexible enough to handle rotational symmetries or axis invariances by removing or repositioning anchor points as needed (see Appendix -C for details on modifying anchor points for a rotationally symmetric object). Apart from the adjustments for rotationally symmetric objects, we use the same value for hyperparameter $L$ for all tasks in the experiments. This highlights the generality and robustness of the proposed reward specification, making it broadly applicable across various tasks and objects.
Figure 4: Object Pose Tracking Reward. The agent is rewarded for moving the object along a goal trajectory by minimizing relative pose distance $d(\bm{T}^{\text{goal}}_{\tau+t},\bm{T}^{\text{obj}}_{t})$, computed using anchor points $\{\bm{k}_{i}\}_{i=1}^{N^{\text{anchor}}}$ in the local frame, where $L$ is a distance parameter for orientation. For the goal trajectory, the $t=\tau$ offset accounts for the start of the demonstration when the object is stationary, which is removed from the goal trajectory for reward computation.
The observation at timestep $t$, $\bm{o}_{t}$ is defined as
$\centering\bm{o}_{t}=[\bm{q}_{t},\bm{\dot{q}}_{t},\bm{X}^{\text{fingertips}}_{t},\bm{x}^{\text{palm}}_{t},\bm{X}^{\text{obj}}_{t},\bm{X}^{\text{goal}}_{t}]\@add@centering$ (3)
where $\bm{q}_{t}\in\mathbb{R}^{N^{\text{joints}}}$ is the vector of robot joint angles, $\bm{\dot{q}}_{t}\in\mathbb{R}^{N^{\text{joints}}}$ is the vector of robot joint velocities, $\bm{X}^{\text{fingertips}}_{t}\in\mathbb{R}^{N^{\text{fingers}}\times 3}$ contains 3D positions of each fingertip, $\bm{x}^{\text{palm}}_{t}\in\mathbb{R}^{3}$ is the palm center’s 3D position, and $\bm{X}^{\text{obj}}_{t},\bm{X}^{\text{goal}}_{t}\in\mathbb{R}^{N^{\text{anchor}}\times 3}$ contain 3D positions of the anchor points representing the current and goal object pose, respectively. $N^{\text{joints}},N^{\text{fingers}}\in\mathbb{N}$ are the number of robot joints and fingers, respectively.
The action at timestep $t$, $\bm{a}_{t}$, is defined as
$\centering\bm{a}_{t}=[\bm{x}^{\text{palm-target}}_{t},\bm{r}^{\text{palm-target}}_{t},\bm{x}^{\text{pca-target}}_{t}]\@add@centering$ (4)
where $\bm{x}^{\text{palm-target}}_{t}\in\mathbb{R}^{3}$ is the target palm center position, $\bm{r}^{\text{palm-target}}_{t}\in\mathbb{R}^{3}$ represents the target palm orientation expressed as Euler angles, and $\bm{x}^{\text{pca-target}}_{t}\in\mathbb{R}^{N_{\text{pca}}}$ represents the target PCA values used to control the hand configuration, where $N_{\text{pca}}=5$ is the number of principal components. We employ a geometric fabric controller and action space similar to the approach introduced in Lum et al. [30]. In this framework, hand configurations are controlled via a lower-dimensional PCA representation, enabling human-like grasping and manipulation. For more details on this action space formulation and its benefits, we refer readers to their work.
We use an initial state distribution, guided by the human pre-manipulation hand pose, to simplify exploration and bias the policy toward learning human-like behavior. We construct this distribution by sampling the object pose around the trajectory’s initial object pose, then set the robot configuration to match the pre-manipulation hand pose with slight perturbation (see Appendix -E for details). Given this perturbed hand pose, we compute a feasible arm joint angle configuration with IK  [^1] We use cuRobo [48] to perform parallelized, collision-free IK to ensure RL training performance is not bottlenecked by IK computations. The environment is reset if the object is too far from the goal $d(\bm{T}^{\text{goal}}_{t},\bm{T}^{\text{obj}}_{t})>D_{\text{max}}$, the robot palm is too far from the object $d(\bm{T}^{\text{palm}}_{t},\bm{T}^{\text{obj}}_{t})>D_{\text{max}}$, or the total timesteps of the goal trajectory is reached $\tau+t>T$, where $D_{\text{max}}=0.25m$.
When capturing the real-world scene, the exact dynamics parameters (e.g., the object’s mass, inertia, and friction, the friction properties of the table and robot hand) are often unknown. To address this uncertainty, we employ domain randomization during simulation training, enabling the resultant policy to perform adaptively in a diverse range of dynamic environmental conditions and facilitating sim-to-real transfer.
We train an LSTM-based policy capable of leveraging a history of observations to extract temporal information and address challenges such as noisy or incomplete data, hidden states, and partial observability. By incorporating sequential context, the policy can intelligently handle scenarios involving noisy object poses, unknown system dynamics parameters, and other sources of uncertainty to produce more effective actions. We train the policy with object pose instead of image observations, speeding up training due to the low-dimensional observation space and the absence of image rendering, while also enabling robustness to visual changes and distractors. We also train the policy with object pose observation noise to improve its robustness to pose prediction errors and camera calibration errors. Additionally, we introduce object perturbations using random forces. This, together with extensive domain randomization, improves the policy’s ability to policy handle unexpected contacts, physical disturbances, and variations in dynamic parameters, resulting in robust zero-shot sim-to-real policy transfer. See Appendix -F for more training details.
III-C Sim-to-Real
Figure 5: Inference-Time Diagram. The RL policy takes object pose and robot proprioception as input. It outputs an action that is sent to a geometric fabric controller [30], which generates joint PD targets. See more at Appendix -D.
Figure 6: Hardware Setup & Digital Twin. The hardware setup includes an Allegro hand mounted on a KUKA LBR iiwa 14 arm with a ZED 1 stereo camera mounted on the table (images from the camera’s perspective). Experiments are conducted in a tabletop setting with a static box, saucepan, and dishrack, involving manipulation tasks with three objects: snackbox, pitcher, and plate. The right image illustrates our digital twin.
After training, we deploy the policy on a real robot without additional fine-tuning (i.e., zero-shot). Figure 5 illustrates the control pipeline at deployment, highlighting the inputs and outputs of our policy at test time. We track object 6D poses at 30Hz using FoundationPose [54]. The RL policy processes real-time observations and outputs actions at 15Hz, which are then passed to a geometric fabric controller [30] running at 60Hz. Finally, this controller produces robot joint PD targets, which are executed by a low-level PD controller at 200Hz.
IV Experiments & Results
Our experiments aim to answer the following questions:
  1. 1.
    Importance of Embodiment-Specific RL: Do RL policies trained via Human2Sim2Robot outperform baselines on dexterous manipulation tasks?
  2. 2.
    Importance of Object Pose Trajectory: How effective is the object pose trajectory from a human demonstration as a dense reward for RL policy training, compared to other reward formulations?
  3. 3.
    Importance of Pre-Manipulation Pose Initialization: Does a pre-manipulation hand pose from a human demonstration provide more effective initialization for learning manipulation skills than generic initializations?
  4. 4.
    Sufficiency of Pre-Manipulation Hand Pose: How effective is a single pre-manipulation pose in guiding RL policy training, compared to alternatives that require full human hand trajectories?
We evaluate Human2Sim2Robot in simulation and on a real robot across a diverse set of tasks and objects to answer these questions, which we describe in more detail below.
IV-A Experimental Setup
Hardware Setup. Figure 6 shows our real-world hardware setup and objects, as well as our corresponding digital twin in simulation. Our robot consists of a 16-DoF dexterous Allegro hand mounted on a 7-DoF KUKA LBR iiwa 14 arm. We use a ZED 1 stereo camera mounted to the table for both recording the human demonstration and real-time object pose estimation for policy input at test time. Our experiments are conducted in a tabletop setting with three static objects on the table: a box, a large saucepan placed atop the box, and a dishrack. The tabletop and its static objects are captured in scene scan $\mathcal{S}$.
Figure 7: Task Visualization. We visualize our real-world tasks spanning grasping, non-prehensile manipulation, and extrinsic manipulation, with the human demonstration and corresponding real robot behavior. We also have 3 complex multi-step tasks composing multiple skill types shown above: snackbox push-pivot, plate lift-rack, plate pivot-lift-rack.
Tasks & Objects. We perform experiments with three objects: snackbox, pitcher, and plate. Figure 7 visualizes our tasks using these objects, which include grasping, non-prehensile manipulation, and extrinsic manipulation. We also explore multi-step tasks that compose sequences of these skills, such as pivoting the plate, lifting it, and then placing it in a dishrack (see Appendix -G for the full task list).
IV-B Importance of Embodiment-Specific RL
To highlight the benefits of a closed-loop RL policy, we compare Human2Sim2Robot to baselines that do not learn policies via RL. Without RL, these baselines require robot action labels for the entire task, which is achieved by performing human hand pose estimation and human-to-robot retargeting for every frame of the video demonstration. See Appendix -H for details on baselines. We evaluate Human2Sim2Robot against 3 baselines across 7 real-world tasks:
  • Replay: Performs open-loop replay of the robot joint configuration trajectory by setting the joint PD targets to the retargeted joint positions.
  • Object-Aware (OA) Replay: Like Replay, but warps the trajectory by the relative transform between initial object pose in the human demo and at test time (similar to PEP [51] and OKAMI [27]).
  • Behavior Cloning (BC): Trains a closed-loop diffusion policy [8] on 30 demos (same number as HuDOR’s [16] baseline), generated from our one demo by sampling object poses (same range as our RL training), performing OA Replay, and using these trajectories as demo data.
Figure 8: Real-World Success Rates. Human2Sim2Robot’s RL policies outperform Replay by 67%, Object-Aware (OA) Replay by 55%, and Behavior Cloning (BC) by 68%, across grasping, non-prehensile manipulation, and complex tasks composing multiple skill types.
Figure 9: Plate Pivot Lift Rack. Robot converges on a strategy that is guided by the human demonstration, but adapted to its morphological differences.
For each real-world task, we evaluate the success of the task-specific RL policy and the corresponding baselines on 10 policy rollouts. The results are visualized in Figure 8. Across all tasks, Human2Sim2Robot policies outperforms the baselines substantially. As we hypothesized, Replay’s open-loop trajectory is not successful on most tasks. OA Replay accounts for randomizations in initial position, but still had many failures to due to (1) hand pose estimation errors that make retargeting challenging, (2) morphological differences between the human and Allegro hand, and (3) non-reactive open-loop control. It does perform well on tasks that do not require high precision like snackbox-pivot. BC performed similarly to Replay, which can be attributed to challenges learning from a small, lower quality dataset, as actions are computed from noisy human hand pose estimations. Despite being a closed-loop policy, small deviations from the data distribution resulted in compounding errors throughout policy rollouts. While retargeted robot demos can succeed on simpler tasks, they often fail on harder multi-step plate tasks. In contrast, Human2Sim2Robot policies do not simply imitate the human behavior, but adapt the behavior for the robot embodiment to perform the task, resulting in much higher success rates across tasks studied in our experiments.
Regarding Human2Sim2Robot failure modes, failures typically arose from converging on policies that exploited simulation inaccuracies or from significant pose estimation error from occlusion. Examples of simulation inaccuracies include imperfect friction modeling of the tabletop and static objects, such that policies converged on behavior that leveraged these inaccurate parameters. The most challenging task was plate-pivot-lift-rack (Figure 9) as it required high precision object handling: the thin plate is hard to manipulate and easily slips out of the large Allegro hand.
IV-C Importance of Object Pose Trajectory
We perform experiments in simulation to study the impact of the object pose trajectory-tracking reward. In particular, we study the speed and stability of learning, the final achieved reward, and the qualitative behavior. We ablate this component of our framework by comparing with the following:
  • Fixed Target: In $r^{\text{obj}}_{t}$ (Equation 1), we replace current goal object pose $\bm{T}^{\text{goal}}_{\tau+t}$ with final object pose $\bm{T}^{\text{goal}}_{T}$.
[…]
  • Downsampled Trajectory: In $r^{\text{obj}}_{t}$ (Equation 1), we replace the current goal object pose $\bm{T}^{\text{goal}}_{\tau+t}$ with the downsampled pose $\bm{T}^{\text{goal}}_{\tau+t_{\text{down}}}$, where $t_{\text{down}}=\lfloor t/D\rfloor\cdot D$ and $D$ is the downsampling factor. This reduces the temporal resolution of the trajectory into a series of waypoints.
Figure 10: Object Pose Tracking Reward Ablation. Reward curves comparing our approach with three approaches ablating the object pose tracking reward.
In Figure 10, we report the reward curves of the different approaches on the plate-pivot-lift-rack task. We train policies with three random seeds for all approaches. We see that Human2Sim2Robot achieves a substantially higher average reward than the methods ablating the object pose tracking reward. The plate lying flat on the table is too large to be directly grasped, therefore the optimal strategy is to use extrinsic manipulation leveraging the wall to pivot the plate into a graspable position, as demonstrated in the human video. Fixed Target and Interpolated Target encourage the policy to greedily move the plate directly to the target in the air, but they struggle to pick up the plate and get stuck in a local minimum. The policy does not explore extrinsic manipulation because this requires navigating to low reward regions for a long period. Downsampled Trajectory is able to learn the task, but it takes longer to learn due to the weaker learning signal. It produces jerky motions in hopes of maximizing reward by tracking the waypoints that jump suddenly. Ours uses the full dense object pose trajectory, which ultimately achieves the strongest performance and converges the fastest, underscoring the effectiveness of an object-centric, embodiment-agnostic reward function for learning dexterous manipulation policies with RL that can be successfully deployed in the real-world.
IV-D Importance of Pre-Manipulation Pose Initialization
We perform experiments in simulation to study the impact of pre-manipulation hand pose initialization. In particular, we study the speed and stability of learning, the final achieved reward, and the qualitative behavior. We ablate this component of our framework by comparing with the following:
  • Default Initialization: We initialize the robot configuration at a default rest pose that is not close to the object.
  • Overhead Initialization: We compute a joint configuration with IK setting the robot palm 5cm above the object. We set the hand joint angles to a default open hand pose.
  • Pre-Manipulation Far: We initialize the robot with the pre-manipulation hand pose, but adjust the arm joint angles to move the robot palm further 20cm away.
Figure 11: Pre-Manipulation Hand Pose Ablation. Reward curves comparing our approach with three approaches ablating the pre-manipulation hand pose.
In Figure 11, we report the reward curves of the different approaches on the plate-pivot-lift-rack task. We train policies with three random seeds for all approaches. For the plate-pivot-lift-rack task, we see that Human2Sim2Robot achieves a substantially higher average reward than the methods ablating the pre-manipulation pose initialization. Default Initialization and Pre-Manipulation Far perform the worst due to exploration challenges from starting with the hand too far away from the object. Overhead Initialization performs slightly better because it is initialized closer to the object, but fails to converge on a successful policy because the overhead grasp provides a disadvantageous prior, as it is not the optimal approach for performing the task. Ours ultimately results in the highest rewards by minimizing the exploration challenges and providing an advantageous initialization for faster learning.
To eliminate the need to always initialize the robot hand in close proximity to the object, methods such as collision-free motion planning or straightforward reward formulations like the L2 distance to the pre-manipulation pose have been extensively studied. These techniques can be seamlessly integrated into our system to facilitate navigation from a default rest pose to the pre-manipulation pose. The ablation tests presented here highlight the significant effectiveness of the pre-manipulation pose in guiding the RL policy toward learning human-like behaviors for contact-rich manipulation tasks.
Qualitatively, for tasks like plate-pivot-lift-rack that are more intricate, small differences in the pre-manipulation hand pose can result in very different learned strategies due to the differences in the human and robot morphologies. For instance, while the human hand used the pinky and ring fingers to lift the plate before transitioning to a grasp, the Allegro hand, which is much larger, used its ring finger to pivot the plate and clipped it between the middle and index finger once the plate was off the table (see Figure 9). This further underscores our hypothesis that significant differences in robot morphologies may lead to strategies that are guided by the human motion but ultimately converge on a different strategy that is more suitable for the robot’s embodiment after learning through trial-and-error.
IV-E Sufficiency of Single Pre-Manipulation Hand Pose
We perform experiments in simulation to compare pre-manipulation hand pose initialization to the following methods that require the full human hand trajectory:
  • Hand Tracking Reward: We add a hand trajectory tracking reward term $r^{\text{hand}}_{t}=\exp\left(-\alpha||\bm{X}^{\text{fingertips}}-\bm{X}^{\text{desired-fingertips}}||\right)$ to our reward function to follow the human hand trajectory for the entire episode. Our new reward function is $r_{t}=r^{\text{obj}}_{t}+r^{\text{hand}}_{t}$.
  • Residual Policy: We keep the same object-centric reward function, but now the retargeted robot configuration trajectory is replayed open-loop, and the action space learns delta PD joint target actions, inspired by Chen et al. [6].
Figure 12: Full Hand Trajectory-Tracking Ablation. Reward curves comparing our approach with two approaches that leverage the full hand trajectory.
In Figure 12, we report the reward curves of the different approaches on the plate-pivot-lift-rack task. We train policies with three random seeds for all approaches. We find that Hand Tracking Reward is able to learn an effective policy, but it learns more slowly because it initially focuses on improving its hand-tracking reward, which is not always conducive to policy performance. When trained to convergence, Hand Tracking Reward does not show any substantial improvement over our method, despite requiring additional data and supervision. Residual Policy’s performance is substantially worse because inaccurate hand pose estimation results in a poor base motion that is difficult to learn an effective residual policy for. In contrast, our method is able to effectively learn the task without the full human hand trajectory as additional supervision.
V Limitations & Future Work
In this paper, we evaluate Human2Sim2Robot’s ability to bridge the embodiment gap between human hands and a Kuka arm and Allegro hand. Future work could extend our evaluation to other robot hands or even substantially different arms and grippers (e.g., parallel-jaw grippers). We note that no aspects of Human2Sim2Robot are specific to the embodiment used in our experiments, and thus we expect it will be applicable to other robot embodiments.
In addition, Human2Sim2Robot focuses on tasks specified by the pose trajectory of a single object. Adapting the method to handle multiple objects is possible with our framework, but would require additional modifications. Human2Sim2Robot also assumes a high-quality object tracker (in our case, an object pose estimator) and simulator. Our tasks therefore only feature rigid-body objects and environments, which can be efficiently tracked with existing pose estimators and simulated using existing rigid-body simulators. Extending to articulated or deformable objects would necessitate adapting Human2Sim2Robot to using different approaches for state estimation and simulation modeling.
While using 6D object pose as policy input enhances robustness against visual distractors and simplifies RL training and sim-to-real transfer, our pose estimator has limitations as it struggles with pose ambiguity of symmetric objects and sensor noise from reflective objects. Invariance to pose ambiguity of symmetric objects could be achieved by modifying the anchor points used to determine object-centric rewards, for instance by automatically determining the axis of symmetry given the object mesh [39] and removing the points orthogonal to this axis of rotation, as detailed in our reward computation definition in Section III-B. For reflective objects, these challenges can be addressed by distilling the pose-based policies into image-based policies, similar to prior work [30, 49].
Finally, Human2Sim2Robot currently trains robust single-object, single-task policies. In future work, we can explore training generalist multi-task, multi-object policies that are conditioned on object shape and desired object trajectory. This can be achieved using multi-task RL or teacher-student distillation. Future work can also investigate extensions or modifications to Human2Sim2Robot for bimanual manipulation or whole-body manipulation tasks.
VI Conclusion
We present Human2Sim2Robot, a real-to-sim-to-real RL framework for learning robust dexterous manipulation robot policies from a single human hand RGB-D video demonstration. Our method facilitates training RL policies in simulation for dexterous manipulation by leveraging object pose trajectories for task specification and a pre-manipulation hand pose initialization for efficient policy exploration, eliminating the need for wearables, teleoperation, or extensive data collection efforts typically required in IL. By formalizing tasks through object-centric rewards and a pre-manipulation hand pose, Human2Sim2Robot addresses several key challenges in general dexterous manipulation, including efficient exploration, eliminating reward engineering effort, crossing the human-robot embodiment gap, and robust sim-to-real transfer. Our experiments demonstrate significant improvements over existing methods across grasping, non-prehensile manipulation, and extrinsic manipulation tasks. Overall, Human2Sim2Robot represents a step forward in facilitating scalable, efficient, and robust training of real-world dexterous manipulation policies using sim-to-real RL with minimal human input.
Acknowledgments
We thank the reviewers for their helpful suggestions and feedback. This work is supported by Stanford Human-Centered Artificial Intelligence, the National Science Foundation under Grant Numbers 2153854 and 2342246, and the Natural Sciences and Engineering Research Council of Canada (NSERC) under Award Number 526541680.